On January 1, 2016, Cayce Corporation acquired 100 percent of Simbel Company for consideration transferred with a fair value of $129,600. Cayce is a U.S.-based company headquartered in Buffalo, New York, and Simbel is in Cairo, Egypt. Cayce accounts for its investment in Simbel under the initial value method. Any excess of fair value of consideration transferred over book value is attributable to undervalued land on Simbel’s books. Simbel had no retained earnings at the date of acquisition. Following are the 2017 financial statements for the two operations. Information for Cayce and for Simbel is in U.S. dollars ($) and Egyptian pounds (£E), respectively.

Cayce
Corporation
Simbel
Company
Sales $ 208,000 £E 823,100
Cost of goods sold (97,800 ) (432,100 )
Salary expense (20,000 ) (76,000 )
Rent expense (7,500 ) (47,000 )
Other expenses (22,500 ) (60,500 )
Dividend income—from Simbel 15,125 0
Gain on sale of building, 10/1/17 0 35,000
Net income $ 75,325 £E 242,500
Retained earnings, 1/1/17 $ 323,000 £E 137,000
Net income 75,325 242,500
Dividends (29,000 ) (55,000 )
Retained earnings, 12/31/17 $ 369,325 £E 324,500
Cash and receivables $ 111,300 £E 151,500
Inventory 98,500 308,000
Prepaid expenses 30,000 0
Investment in Simbel (initial value) 129,600 0
Property, plant & equipment (net) 414,000 460,000
Total assets $ 783,400 £E 919,500
Accounts payable $ 62,800 £E 55,500
Notes payable—due in 2020 139,775 141,500
Common stock 125,000 245,000
Additional paid-in capital 86,500 153,000
Retained earnings, 12/31/17 369,325 324,500
Total liabilities and equities $ 783,400 £E 919,500

Additional Information

  • During 2016, the first year of joint operation, Simbel reported income of £E 168,000 earned evenly throughout the year. Simbel declared a dividend of £E 31,000 to Cayce on June 1 of that year. Simbel also declared the 2017 dividend on June 1.

  • On December 9, 2017, Simbel classified a £E 10,500 expenditure as a rent expense, although this payment related to prepayment of rent for the first few months of 2018.

  • The exchange rates for 1 £E are as follows:

January 1, 2016 $ 0.300
June 1, 2016 0.290
Weighted average rate for 2016 0.288
December 31, 2017 0.280
June 1, 2017 0.275
October 1, 2017 0.273
Weighted average rate for 2017 0.274
December 31, 2017 0.270

Translate Simbel’s 2017 financial statements into U.S. dollars and prepare a consolidation worksheet for Cayce and its Egyptian subsidiary. Assume that the Egyptian pound is the subsidiary’s functional currency.

req B consolidation worksheet

adj. & eliminations

consolidated

balances

account

cayce

dollars

simbel

dollars

debit credit dollars
sales
COGS
salary expense
rent expense
other expenses
dividend income
gain on sale of fixed asset 10/1/17
net income
retained earnings 1/1/17
net income
dividends
retained earnings 12/31/17
cash & receivables
inventory
prepaid rent
investment
PPE
total assets
a/p
notes payable
common stock
APIC
retained earnings 12/31/17
subtotal
cumulative translation adjustment (positive or negative)
total liabilities & equities

THANKS!
Answer #1

Answer:

Req b

2.Preparation of Consolidated Worksheet for Cayoe Corporation: Comsolidated Workaheet of Cayce Corporation Cayoe ($5 Simbel (2|Dobir S Crodit 3 Paleace (9 Accounts 208,000 225,529 (118,395) (20,000(20,824) (7,500)(10,001) (22,500)16,577) 433,529 CIS1) (40,824) (17,501) Sales 390 Other expense Dividend income Gain on sale of buiding, 101/17 Net Income 15,125 152 9.555 75,325 69,287 179.958 323,00039,394 39,394 39,394 362,394 Retained carnings 1/1/17 Net Income 75,32569,287 15,125 (凭2.MI) 447,756 Retained earningg 12/31/17 Cash and receivables Inventory Investment in Simbel initial valuc 152,205 181,660 30,000 129,600 2,835 414.000 124.200 Totalasset783,400 251,100 62,80014,985 139,77538,205 t & net 914,080 Accounts payable Notes payable-due in 2020 Common stock Additional paid-in capital Retained carnings, 12/31/17 Subtotal Cumulative trans lation adjustment 177,980 125,00073,500 73,500 86,500 45,90045,900 462,881 930,146 16 1100 914,080 369,32593,556 266,146 Total Liabilities and equity783,400 251,100

Working Notes :

Working Notes :

1. Prepayment of rent made for 2018 for £E 10,500 should be reclassified from rent expense to prepaid rent paid in Simbel's book.

2. Translation of Retained earnings on 1/1/17-

Retained earnings 1/1/16 -0- -0-
Net income for 2016 (168,000) 0.288 (48,384)
Dividend paid 6/1/16 31,000 0.290 8,990
Retained earnings, 1/1/17 £E(137,000) $(39,394)

3. Calculation of cumulative translation adjustment:

Pounds

(£E)

Rate

Dollar

($)

Net assets 1/1/16 (245,000 + 153,000) (398,000) 0.300 (119,400)
Net Income, 2016 (168,000) 0.288 (48,384)
Dividends , 6/1/16 31,000 0.290 8,990
Net assets, 12/31/16 (535,000) (158,794)
Net assets, 12/31/16 at current exchange rates (535,000) 0.280 (149,800)
Translation adjustment, 2017 (negative) (1) (8,994)
Net assets, 1/1/17 (535,000) 0.280 (149,800)
Net Income, 2017 (253,000) from translation sheet (69,287)
Dividends, 6/1/17 55,000 0.275 15,125
Net assets , 12/31/17 (733,000) (203,962)
Net assets 12/31/17 at current exchange rate (733,000) 0.270 (197,910)
Translation adjustment, 2017 (negative) (2) (6,052)
Cumulative translation adjustment, 12/31/2017 (negative) (1) + (2) $(15,046)

4. Calculation of excess of cost over book value -

Purchase Price (given) $129,600
Less :

Common stock

(73,500)

Additional paid in capital

(45,900)
Excess of purchase price over book value $10,200

5. Excess of cost over book value is = ( 10,200 ÷ 0.300) = $34,000

Equivalent US dollar at 12/31/18 = 34,000 * 0.270 = $ 9,180

Equivalent US dollar at 1/1/17 = 34,000 * 0.300 = $ 10,200

Cumulative translation adjustment (negative) = $9,180 - 10,200 = $ 1,020